An electrical space heater contains a Nichrome heating element that has a resistance
of 8.00 Ω at 20.0 C. When 120 V are applied across the heating element, the electric
current heats the Nichrome wire to 1000C.
a) Show the complete calculation of the heating element resistance at 1000 C.
b) Show the complete calculation of the heating element current at 1000 C.
c) Show the complete calculation of the heating element operating power at 1000 C.
1
Expert's answer
2020-02-10T09:46:26-0500

Solution. a) Resistance values for conductors at any temperature must be determined through yet another formula:


"R_t=R_0(1+\alpha(T-T_0))"

where Rt is resistance for temperature T=10000C; T0 =200C is initial temperature; α=0.00017 C-1 temperature coefficient of resistance; R0=8Ω is resistance for temperature T=200C. Therefore get


"R_t=8(1+0.00017(1000-20))=9.33\Omega"

b) Using Ohm's law we get


"I=\frac{U}{R} = \frac {120V}{9.33\Omega}=12.86A"

c) Using the electric power formula, we obtain


"P=UI=120\times12.86=1543.2W"

Answer. a) Rt=9.33Ω; b) I=12.86A; c) 1543.2W